OpenStudy (anonymous):

what is the square root of 150

OpenStudy (adll23):

12.2474487

OpenStudy (anonymous):

they don't have that

OpenStudy (adll23):

so if it was rounded it would be 12.25

OpenStudy (adll23):

what are the choices?

OpenStudy (anonymous):

@adll23 no direct answers

OpenStudy (trojanpoem):

sqrt(150) = sqrt( 3 * 50) = sqrt(6 * 25) = 5sqrt(6)

OpenStudy (adll23):

option A

OpenStudy (anonymous):

c

OpenStudy (adll23):

150 = 2 * 3 * 5 * 5, so 5sqrt(6)

OpenStudy (anonymous):

wait wat is it a or c

OpenStudy (trojanpoem):

A

OpenStudy (anonymous):

my mistak a

OpenStudy (anonymous):

what is square root of 200

OpenStudy (dayakar):

\[\sqrt{25*6}\]

OpenStudy (adll23):

√ 200 = √ [ 2 ( 10 ) ² ] = ± 10 √ 2

OpenStudy (anonymous):

what is the square root 169 + square root 50

OpenStudy (anonymous):

what is the square root of 169 + square root of 81

OpenStudy (whovianchick):

square root 169= 13 square root 81= 9 added=22

OpenStudy (whovianchick):

square root of 50= 5squareroot2 or 7.07 added to 13= 20.07

OpenStudy (anonymous):

Simplify the expression what is the square root of 18+ square root of 50

imqwerty (imqwerty):

hello i jst came here to smile :)

OpenStudy (alexandervonhumboldt2):

I came here to watch everything happening here. Please post only 1 question per post. That is because people will get medals and it will raise your smartscore.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

Simplify the expression what is the square root of 18+ square root of 50

OpenStudy (whovianchick):

post a new question so it's fair, don't keep commenting on this one

OpenStudy (anonymous):

wat is the answer

imqwerty (imqwerty):

the answer can be calculated but its a way boring proccess

imqwerty (imqwerty):

can i use a calcultaor?

OpenStudy (alexandervonhumboldt2):

Here are all the answers to your test. I took this earlier, have fun! And oh, I got a 100% so it's totally worth copying off of here: http://pages.citebite.com/a4o3b6r4a2tjy

OpenStudy (alexandervonhumboldt2):

please post only 1 question per post, this is the rule.

OpenStudy (trojanpoem):

@AlexandervonHumboldt2 , TROLOLO

OpenStudy (alexandervonhumboldt2):

openstudy is made for studying not for throwing away answers.

OpenStudy (alexandervonhumboldt2):

well, this is actually made by @♪Chibiterasu , i just use it

OpenStudy (anonymous):

yall is no help lol

OpenStudy (trojanpoem):

@AlexandervonHumboldt2 , I thought it was really her test answer. 1 medal for that.

OpenStudy (alexandervonhumboldt2):

well it is, i posted a answer to her request of answer: my answer is read the anti-cheating policy girl

OpenStudy (alexandervonhumboldt2):

OMG 3 medals for a joke

OpenStudy (trojanpoem):

@AlexandervonHumboldt2 , You deserve more.

OpenStudy (alexandervonhumboldt2):

maybe xD

OpenStudy (alexandervonhumboldt2):

did you passed your test? @tasha378

OpenStudy (alexandervonhumboldt2):

have you learned anything? how to simplify sqrt150? oh yeah you havent learned, you just wanna 100000000000%

OpenStudy (anonymous):

1. Find the largest perfect square that divides evenly into 150. That number is 25. 2. Rewrite sqrt (150) = sqrt (25) * sqrt (6) 3. Square root of 25 is 5 so replace and get 4. 5 * sqrt (6)

OpenStudy (anonymous):

12.25